• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name E.Chudakov

    Log entry time 00:06:59 on April 5,2008

    Entry number 229929

    keyword=Shift summary

    16:00 MCC is retuning the beam
    16:45 Beam is back at 5uA
    HV tests:
    BB DC HV=1500V: runs 2294-2295
    BB DC HV=1550V: runs 2296
    19:00 MCC stopped the beam in order to tune HallC
    20:30 MCC starts to restore the beam.

    20:30 Trying to run with BB magnet off: all the HV trip...

    21:50 We gave up on trying to run with BB magnet off
    BB set to 390A
    HRSL collimator/sieve removed
    HV DC is left at 1550V

    Running at 4.8uA, no sieve
    22:00 run 2298 0.46M events, stopped when DC tripped and ROC5 died
    22:20 run 2299 0.22M: stopped to reboot ROC5
    22:30-23:20 runs 2300-2301: 0.4M

    Running at 2.7uA
    2300-2304